A particle of mass 1 kg is projected with velocity 20√2m/s. at45° with ground when the particle is at a highest point angular velocity about point of projection

Grade:12th pass

1 Answers

jitender
6 years ago
V at high as point V = 20√2cos45= 20m/sDisplacement from initial point to highest point D = √(H^2+(R/2)^2) Where H =[( 20√2) ^2 sin^2 45]/2g = 40mAnd R=[( 20√2) ^2 sin 2×45]/g = 80mNow D = 40√2mAngular velocity ₩ = V_/DV_= velocity perpendicular to DV_ = Vcos 45 = component of V perpendicular to D₩ = (20×1/√2)/40√2 = 1/4 = 0.25rad/s